OpenStudy (en):

PLEASE HELP! find the derivative of w=sin^4ycos^4y. my final answer is 4sin^3cos^3y * cos2y. but the answer written in my book is 1/2sin^3 (2y)cos(2y)

OpenStudy (freckles):

I recommend using the double angle identity first to rewrite w then differentiate

OpenStudy (freckles):

\[\sin(2y)=2 \sin(y) \cos(y) \\ \frac{1}{2} \sin(2y)=\sin(y) \cos(y) \\ w=( \frac{1}{2} \sin(2y))^4 \]

OpenStudy (en):

w = sin^4y cos^4y dw/dy = 4 sin^3y* cos^5y - 4 cos^3y* sin^5y = 4 sin^3y cos^3y[ cos^2y - sin^2y] = 4 sin^3y cos^3y * cos2y = 1/2*(2siny cosy)^3 * cos2y = 1/2 sin^3(2y) cos2y someone answere it this way... i just dont get where the 1/2 came from.. please explain

OpenStudy (freckles):

\[2\sin(y)\cos(y)=\sin(2y)\]

OpenStudy (freckles):

oh you used that

OpenStudy (anonymous):

@En dont copy and past from yahoo answers.. its best to solve them yourself for you can learn more.

OpenStudy (freckles):

\[4=\frac{1}{2}(2)^3 \text{ they just rewrote 4 }\]

OpenStudy (en):

@freckles tnx!

OpenStudy (freckles):

np i was kind of confused at first i thought it was your solution

OpenStudy (freckles):

@En I feel like that first way I suggested might be a little easier

OpenStudy (en):

i solved the 4sin^3cos^3y * cos2y i just got confused in the simplifying process :)

OpenStudy (freckles):

\[w=(\frac{1}{2} \sin(2y))^4 \\ w'=4(\frac{1}{2}\sin(2y))^{3}(\frac{1}{2} \sin(2y))' \\ w'=4(\frac{1}{2}\sin(2y))^3 \frac{1}{2} \cdot 2 \cos(2y) \\ w'=4 (\frac{1}{2})^3 \sin^3(2y) \cos(2y) \\ w'=\frac{4}{8} \sin^3(2y) \cos(2y) \\ w'=\frac{1}{2} \sin^3(2y) \cos(2y)\] just a whole bunch of chain rule

OpenStudy (en):

Where did you get the w=1/2sin(2y))^4? @freckles

OpenStudy (freckles):

the double angle identity 2sin(y)cos(y)=sin(2y) so sin(y)cos(y)=1/2*sin(2y)

OpenStudy (freckles):

so (sin(y) cos(y))^4=(1/2*sin(2y))^4

OpenStudy (freckles):

and (sin(y) cos(y))^4=sin^4(y)*cos^4(y)

OpenStudy (en):

where did the "1/2" came from?

OpenStudy (freckles):

2sin(y)cos(y) is equal to sin(2y) do you understand this part?

OpenStudy (freckles):

divide both sides by 2 sin(y)cos(y) is equal to 1/2*sin(2y)

OpenStudy (en):

yes